OpenStudy (oleg3321):

please help with algebra 1

OpenStudy (oleg3321):

A system of equations is given below. 2x + 7y = 1 -3x – 4y = 5 Create an equivalent system of equations by replacing the first equation by multiplying the first equation by an integer other than 1, and adding it to the second equation. Use any method to solve the equivalent system of equations (the new first equation with the original second equation). Prove that the solution for the equivalent system is the same as the solution for the original system of equations.